dy/dx= −(x^2/y^2)+1/[3y^2(x3+ y3)^2]


1
Expert's answer
2021-04-13T23:13:05-0400

"\displaystyle
\frac{\mathrm{d}y}{\mathrm{d}x} = -\frac{x^2}{y^2} + \frac{1}{3y^2(x^3 + y^3)^2}\\

3y^2(x^3 + y^3)^2\frac{\mathrm{d}y}{\mathrm{d}x} = -3x^2(x^3 + y^3)^2 + 1 \\

3y^2(x^3 + y^3)^2\frac{\mathrm{d}y}{\mathrm{d}x} + 3x^2(x^3 + y^3)^2 = 1\\

3(x^3 + y^3)^2\left(y^2\frac{\mathrm{d}y}{\mathrm{d}x} + x^2\right) = 1\\

3(x^3 + y^3)^2\left(y^2\frac{\mathrm{d}y}{\mathrm{d}x} + x^2\right) = 1\\

\frac{1}{3}\frac{\mathrm{d}}{\mathrm{d}x}\left(x^3 + y^3\right)^3 = 3\\

\left(x^3 + y^3\right)^3 = \int \mathrm{d}x = 3(x + C)\\

x^3 + y^3 = \sqrt[3]{3(x + C)}\\

y^3 = \sqrt[3]{3(x + C)} - x^3\\

y= \sqrt[3]{\sqrt[3]{3(x + C)} - x^3}"